OpenStudy (seb.cal):

The graph of f(x) = 2x + 1 is shown below. Explain how to find the average rate of change between x = 0 and x = 3.

OpenStudy (seb.cal):

The 'x' is raised as an exponent.

OpenStudy (seb.cal):

@welshfella @ILovePuppiesLol help

OpenStudy (seb.cal):

@imqwerty @Mehek14

OpenStudy (seb.cal):

@nincompoop @sleepyjess please help.

OpenStudy (.mehek.):

Put 0 for x and find what f(0) is equal to Do the same for x = 3

OpenStudy (seb.cal):

ok, and that'll be the answer..? or is that just one step?

OpenStudy (.mehek.):

f(0) = 2 * 0 + 1 = ? f(3) = 2 * 3 + 1 = ?

OpenStudy (.mehek.):

First step

OpenStudy (seb.cal):

f(0) = 1 and f(3) = 7

OpenStudy (.mehek.):

Ok now we use the formula to find rate of change \(\dfrac{y_2-y_1}{x_2-x_1}\) \(\dfrac{7-1}{3-0}\)

OpenStudy (seb.cal):

= 6/3 --> 2!

OpenStudy (.mehek.):

There you go :)

OpenStudy (seb.cal):

Thanks! I was getting caught up in the formula process. I was putting in 9 instead of 7 :P

OpenStudy (seb.cal):

Thank you :)

OpenStudy (.mehek.):

You're welcome ^_^